# The Esethu Framework: Reimagining Sustainable Dataset Governance and Curation for Low-Resource Languages

### v1

@@ Abstract
This paper presents the Esethu Framework, a sustainable data curation framework specifically designed to empower local communities and ensure equitable benefit-sharing from their linguistic resources. This framework is supported by the Esethu license, a novel community-centric data license. As a proof of concept, we introduce the Vuk’uzenzele isiXhosa Speech Dataset (ViXSD), an open-source corpus developed under the Esethu Framework and License. The dataset, containing read speech from native isiXhosa speakers enriched with demographic and linguistic metadata, demonstrates how community-driven licensing and curation principles can bridge resource gaps in automatic speech recognition (ASR) for African languages while safeguarding the interests of data creators. We describe the framework guiding dataset development, outline the Esethu license provisions, present the methodology for ViXSD, and present ASR experiments validating ViXSD’s usability in building and refining voice-driven applications for isiXhosa.
@@ 1 Introduction
In this paper, we present the Esethu Framework, a pioneering data governance approach that fundamentally reimagines how low-resource language datasets are created, curated, and sustained. At its core, the framework introduces a novel economic model where licensing revenue is systematically reinvested into dataset expansion, ensuring continuous growth while directly benefiting the communities that create the data. This is supported by the Esethu License, a first-of-its-kind community-centric licensing scheme that maintains community agency over data assets while enabling broad research access and creating clear pathways for ethical commercialization.
As a proof-of-concept, we develop and release the Vuk’uzenzele isiXhosa Speech Dataset (ViXSD). This dataset seeks to build inclusivity for the isiXhosa people and features diverse speakers, covering different genders, accents and backgrounds while demonstrating a crucial sustainable framework to ensure its growth. We detail the data creation process and framework as well as licensing with the aim of empowering researchers and developers to create and refine ASR datasets and models.
We focus on isiXhosa, a Bantu language spoken by more than 9 million people in South Africa, as it epitomizes the low-resource gap Roux and Visagie (2007). As shown in Figure 1, isiXhosa is notably underserved, only providing approximately 61 hours of speech data despite its sizable population, which positions it below the comparative trend line for African languages. Moreover, existing datasets for under-resourced languages rarely offer continual economic benefits to the communities that supply the data. This situation highlights the importance of holistically addressing both data scarcity and community empowerment (Le Ferrand et al., 2022; Nekoto et al., 2022, 2020).
Language Population Region Dialects Phonology Morphology Sounds
[…]
Table 1: Characteristics of the isiXhosa Language
@@ 2 Related Work
The growing demand for speech data sets in underrepresented languages has driven various resource creation efforts. This section reviews existing community-centered initiatives and African speech datasets.
Multiple projects have aimed to address resource gaps in African languages through community involvement. Nekoto et al. introduced the Oshiwambo data set project, which created more than 5000 sentences in the Oshindonga dialect along with their English translations. This participatory and cost-effective approach highlighted the potential for collaborative resource creation in underrepresented African languages, but they were not able to release the dataset due to the lack of an African-centered data license that would benefit the data creators Aubra Anthony, Lakshmee Sharma, and Elina Noor, (2024). Similarly, AfroDigit presented the first audio digit dataset in 38 African languages enabling speech recognition applications such as telephone and street number recognition (Emezue et al., 2023). BibleTTS provides 86 hours of high-fidelity Bible recordings in 10 African languages (Meyer et al., 2022), and AfricanVoices focuses on collecting high-quality speech datasets for African languages while also providing speech synthesizers (Ogun et al., 2024). AfriSpeech-200 extends these efforts by providing a 200-hour pan-African speech corpus for English-accented ASR in the clinical and general domains  (Olatunji et al., 2023b). Meanwhile, Kencorpus features 5.6 million words and 177 hours of speech of 3 Kenyan languages; Swahili, Dholuo, and Luhya (Wanjawa et al., 2022).
Several datasets have been developed to increase ASR resources for isiXhosa. Louw et al. (2001) collected a spontaneous monolingual corpus of isiXhosa and five other South African languages. This data, sourced from first-language speakers via an annotated telephone-based database. Although this work demonstrated the potential of African language resources, its scale and scope were limited. Ström (2018) captured linguistic diversity of isiXhosa by emphasizing the importance of including regional accents and tonal variations. Lastly, Van Der Westhuizen and Niesler (2017) curated a multilingual dataset of four South African languages, including isiXhosa, in a code-mixed setting with English. This resource advanced acoustic modeling and multilingual ASR systems Biswas et al. , demonstrating the potential of code-mixed datasets. However, its scope remained task-specific, limiting its utility for downstream applications.
@@ 3 The Esethu Framework
Figure 2: Sustainable community-driven dataset curation framework
@@ 3.1 Licensing
Inspired by initiatives such as the Nwulite Obodo Data Licenses Chijioke Okorie, Melisa Omino, Christiaan Lombard, Caroline Muchiri Wanjiru, Takudzwa Matekwe, Vukosi Marivate, (2024) and the Kaitiakitanga Māori Data Sovereignty Licenses Taiuru, Karaitiana , we have developed a novel license for this dataset. This license consists of two components: a commercial license and an open license. In essence, the license does not restrict research use and permits commercial use by African entities. However, non-African commercial entities are required to pay a licensing fee, see Section A.1.
Proceeds from the dataset are legally mandated to be reinvested in the creation of more data for the language in question through a local data creation partner, see Section A.2. This ensures a circular system that not only promotes the generation of additional data, benefiting the language community in question through language technology opportunities, but also guarantees that the revenue generated from the dataset supports fairly paid work for speakers. By ensuring that the data is accessible to African commercial entities only, we aim to foster innovation within the African language ecosystem while preserving opportunities for research on the language.
@@ 4 Vuk’uzenzele isiXhosa Speech Dataset (ViXSD)
As a proof of concept, we develop and release the Vuk’uzenzele isiXhosa Speech Dataset (ViXSD), developed with the Esethu framework in mind, and release it under the Esethu License. This section describes the details the language itself, the dataset curation and creation
@@ 4.5 Recording Process and Quality Control
Quality Assurance Measures: To ensure recording quality, participants were first required to submit sample recordings. These were evaluated for background noise, reading fluency, and pacing. Based on this assessment, individualised feedback and recommendations for improvement were provided via email. For instance, some participants defaulted to English when reading numbers. These participants were instructed to read numbers in isiXhosa, such that at least half of the dataset contains numerals expressed in isiXhosa rather than in English. This approach helps preserve linguistic features that may be diminishing in modern speech due to language shift and increasing code-switching trends.
@@
Table 2: Characteristics of the isiXhosa ASR data
@@ 4.6 Dataset
Table 3: Geographic Distribution of Speakers
Data Quality: Participants recorded the dataset in stereo using personal devices, typically in non-studio environments. This introduces variances including domestic sounds such as children, animals, and environmental noises. However, these real-world conditions can improve the robustness of the ASR system, helping it generalize better across diverse, noisy environments.
@@ 4.7.1 Results
Figure 4: Image shows the reduction in common ASR errors when finetuned on the ViXSD data
@@ 4.8 Lessons Learned
Transcriptions: Transcripts should be long form and conversational. This facilitates both easier quality control and a smoother workflow. Technical text or acronyms hinder the transcription processes as annotators often need to consult with the project managers to standardize how these are spoken.
@@
Recording Equipment Considerations: Participants employed various devices resulting in audio captured at different sample rates and in stereo formats. While these technical differences required resampling and channel conversion, we observed no substantial quality gap between these recordings. This suggests that a strict device requirement may be unnecessary if proper recording guidelines are followed.
@@ 5 Conclusion
In this paper, we present the Esethu Framework, a pioneering approach to sustainable dataset governance that fundamentally reimagines how low-resource language data is created, licensed, and grown. Through its innovative reinvestment model, where licensing revenue is systematically channeled back into dataset expansion, the framework establishes a self-sustaining ecosystem for continuous resource development. As a proof of concept, we developed the Vuk’uzenzele isiXhosa Speech Dataset (ViXSD), which demonstrates both the framework’s practical viability and its potential impact: our experiments show that fine-tuning yields up to 4.6% improvement in ASR performance, validating the technical quality of datasets produced under this approach. The Esethu Framework, with its emphasis on community ownership, sustainable growth, and ethical commercialization, offers a reproducible model for addressing the critical resource gaps in other low-resource languages, ensuring that language communities remain primary beneficiaries of their data contributions while creating pathways for economic empowerment through language technology development.
@@ 6 Limitations
Given the ViXSD dataset focuses on news type articles, it may exhibit thematic and linguistic biases. Efforts have been made to preserve dialectal diversity and cultural authenticity, yet idiomatic expressions and culturally specific references may require careful interpretation in downstream NLP applications. Researchers should be cognisant of potential stereotypical or skewed linguistic representations when applying models trained on this corpus. Additionally, although efforts were made to ensure diverse speakers, the current dataset as it stands may not fully representative of the full socio-economic and geographical range of isiXhosa speakers. Future expansions using our proposed framework will include speakers from multiple provinces and regions to broaden coverage.
@@ 7 Ethical considerations
The creation of this ViXSD dataset necessitated careful consideration of ethical issues, particularly regarding data privacy, cultural representation, and linguistic bias. The dataset is derived from publicly available material, ensuring compliance with copyright and data protection regulations. Personally identifiable information (PII) is either absent or anonymised to safeguard individual privacy. Additionally, through the Esethu license, the data curators remain stewards of the dataset and any licensing fees for the data are reinvested into the development of the isiXhosa language speaker community. The isiXhosa contributors were paid a fair rate for their contributions.
@@ Appendix A Appendix
A.1 Esethu License
@@ 3. We may permit commercial use of the dataset on the following terms:
A.2 Extracts from the Commercial Agreement

### v2

@@ Abstract
This paper presents the Esethu Framework, a sustainable data curation framework specifically designed to empower local communities and ensure equitable benefit-sharing from their linguistic resource. This framework is supported by the Esethu license, a novel community-centric data license. As a proof of concept, we introduce the Vuk’uzenzele isiXhosa Speech Dataset (ViXSD), an open-source corpus developed under the Esethu Framework and License. The dataset, containing read speech from native isiXhosa speakers enriched with demographic and linguistic metadata, demonstrates how community-driven licensing and curation principles can bridge resource gaps in automatic speech recognition (ASR) for African languages while safeguarding the interests of data creators. We describe the framework guiding dataset development, outline the Esethu license provisions, present the methodology for ViXSD, and present ASR experiments validating ViXSD’s usability in building and refining voice-driven applications for isiXhosa.
@@ 1 Introduction
In this paper, we present the Esethu Framework, a pioneering data governance approach that fundamentally reimagines how low-resource language datasets are created, curated, and sustained. At its core, the framework introduces a novel economic model where licensing revenue is systematically reinvested into dataset expansion, ensuring continuous growth while directly benefiting the communities that create the data. This is supported by the Esethu License, a first-of-its-kind community-centric licensing scheme that maintains community agency over data assets while enabling broad research access and creating clear pathways for ethical commercialization. In particular, indigenous low-resourced African language speakers stand to benefit most immediately.
As a proof-of-concept, we develop and release the Vuk’uzenzele isiXhosa Speech Dataset (ViXSD). This dataset seeks to build inclusivity for the isiXhosa people and features diverse speakers, covering different genders, accents and backgrounds while demonstrating a crucial sustainable framework to ensure its growth. We detail the data creation process and framework as well as licensing with the aim of empowering researchers and developers, particularly in low-resource language settings, to create and refine ASR datasets and models.
We focus on isiXhosa, a Bantu language spoken by more than 9 million people in South Africa, as it epitomizes the low-resource gap Roux and Visagie (2007). As shown in Figure 1, isiXhosa is notably underserved, only providing approximately 61 hours of speech data despite its sizable population, which positions it below the comparative trend line for African languages. Moreover, existing datasets for under-resourced languages rarely offer continual economic benefits to the communities that supply the data. This situation highlights the importance of holistically addressing both data scarcity and community empowerment (Le Ferrand et al., 2022; Nekoto et al., 2022a, 2020).
@@ 2 Related Work
The growing demand for speech datasets in underrepresented languages has driven various resource creation efforts. This section reviews notable community-centered initiatives and African speech datasets, with a critical lens on their limitations regarding sustained community involvement and ethical licensing.
Multiple projects have aimed to address resource gaps in African languages through community involvement. Nekoto et al. (2022b) introduced the Oshiwambo data set project, which created more than 5000 sentences in the Oshindonga dialect along with their English translations. This participatory and cost-effective approach highlighted the potential for collaborative resource creation in underrepresented African languages, but they were not able to release the dataset due to the absence of an African-centered data license that would benefit the data creators Aubra Anthony, Lakshmee Sharma, and Elina Noor, (2024). This highlights the persistent tension between open data dissemination and ethical stewardship of community-generated content. Similarly, AfroDigit presented the first audio digit dataset in 38 African languages enabling speech recognition applications such as telephone and street number recognition (Emezue et al., 2023). BibleTTS provides 86 hours of high-fidelity Bible recordings in 10 African languages (Meyer et al., 2022), and AfricanVoices focuses on collecting high-quality speech datasets for African languages while also providing speech synthesizers (Ogun et al., 2024). AfriSpeech-200 extends these efforts by providing a 200-hour pan-African speech corpus for English-accented ASR in the clinical and general domains  (Olatunji et al., 2023b). Meanwhile, Kencorpus features 5.6 million words and 177 hours of speech of 3 Kenyan languages; Swahili, Dholuo, and Luhya (Wanjawa et al., 2022). These datasets mark significant progress in resource availability, but remain top-down in structure, often lacking mechanisms for community governance and equitable attribution.
Several datasets have been developed to increase ASR resources specific to isiXhosa. Louw et al. (2001) collected a spontaneous monolingual corpus of isiXhosa and five other South African languages. This data, sourced from first-language speakers via an annotated telephone-based database. Although this work demonstrated the potential of African language resources, its scale and scope were limited. Ström (2018) captured linguistic diversity of isiXhosa by emphasizing the importance of including regional accents and tonal variations. Lastly, Van Der Westhuizen and Niesler (2017) curated a multilingual dataset of four South African languages, including isiXhosa, in a code-mixed setting with English. This resource advanced acoustic modeling and multilingual ASR systems Biswas et al. (2020), demonstrating the potential of code-mixed datasets. However, its scope remained task-specific, limiting its utility for downstream applications.
In summary, while these initiatives demonstrate promising approaches to resource creation, they often fall short in embedding ethical, inclusive and sustainable models of community participation. Future work must prioritize not only linguistic coverage but also frameworks for ethical licensing, co-ownership, and long-term benefit sharing with the communities that contribute their languages.
@@ 3 The Esethu Framework
Figure 2: Sustainable community-driven dataset curation framework.
@@ 3.1 Licensing
Language Speaker Population Region Dialects Phonology Morphology Sounds
[…]
Table 1: Characteristics of the isiXhosa Language.
Alternative Governance Models: Community-owned corporations and co-operatives promote ethical commercialization of data, provided that reinvestment mandates exist in their founding documents. However, these models often restrict open access, limiting academic usage unless further provisions are made; or require substantial legal, financial, and administrative overhead to set up and maintain, which may be prohibitive for many low-resource language communities. Beyond cooperatives, other governance models include data trusts, where data is held and managed by a trustee on behalf of a defined group of beneficiaries, in this case, language communities. While they offer a clear fiduciary duty to the community, establishing a data trust involves complex legal frameworks and may lack flexibility for dynamic participation over time.
@@
Inspired by initiatives such as the Nwulite Obodo Data Licenses Chijioke Okorie, Melisa Omino, Christiaan Lombard, Caroline Muchiri Wanjiru, Takudzwa Matekwe, Vukosi Marivate, (2024) and the Kaitiakitanga Māori Data Sovereignty Licenses Taiuru, Karaitiana , we have developed a novel license for this dataset. This license consists of two components: a commercial license and an open license. In essence, the license does not restrict research use (inline with the Creative Commons BY-NC-SA license) and permits commercial use by African entities. We define African entities as those headquartered in Africa or majority-African-owned; case-by-case waivers permit diaspora or off-continent organizations with demonstrable ties. Non-African commercial entities are required to pay a licensing fee, see Section A.1. We highlight legal strengths and weaknesses of Esethu license in Section A.2.
Proceeds from the dataset are legally mandated to be reinvested in the creation of more data for the language in question, or less resourced languages, through a local data creation partner, see Section A.3. This ensures a circular system that not only promotes the generation of additional data, benefiting the language community in question through language technology opportunities, but also guarantees that the revenue generated from the dataset supports fairly paid work for speakers. By ensuring that the data is accessible to African commercial entities only, we aim to foster innovation within the African language ecosystem while preserving opportunities for research on the language.
@@ 4 Vuk’uzenzele isiXhosa Speech Dataset (ViXSD)
As a proof of concept, we develop and release the Vuk’uzenzele isiXhosa Speech Dataset (ViXSD), developed with the Esethu framework in mind, and release it under the Esethu License with the help of a data curation partner. The data partner handles licensing administration, including collecting and allocating commercial fees. These fees would be reinvested into community-oriented initiatives, supporting continued data collection, compensating contributors, and expanding the dataset. This section describes the details the language itself, the dataset curation and creation.
@@ 4.2 Vuk’uzenzele South African Multilingual Corpus
11footnotetext: https://southafrica-info.com/arts-culture/11-languages-south-africa/22footnotetext: https://glottolog.org/resource/languoid/id/xhos123933footnotetext: such as those found in English
@@ 4.5 Recording Process and Quality Control
Quality Assurance Measures: To ensure recording quality, participants were first required to submit sample recordings. These were evaluated for background noise, reading fluency, and pacing. Based on this assessment, individualised feedback and recommendations for improvement were provided via email. For instance, some participants defaulted to English when reading numbers. These participants were instructed to read numbers in isiXhosa, such that at least half of the dataset contains numerals expressed in isiXhosa rather than in English. This approach helps preserve linguistic features that may be diminishing in modern speech due to language shift and increasing code-switching trends. Also, participants were instructed to read each prompt at a comfortable pace in a quiet room. To ensure high-quality audio, participants were permitted to re-record any prompt if they misread, hesitated, or paused unnaturally. We then ran an automated text–audio alignment tool and manually spot-checked all flagged utterances for major errors. Only those recordings that were clear, intelligible, and closely matched the target text were retained in the dataset. The full list of speaker guidelines is provided in Appendix A.4.
@@
Table 2: Characteristics of the isiXhosa ASR data.
@@ 4.6 Dataset
Table 3: Geographic Distribution of Speakers.
Data Quality: Participants recorded the dataset in stereo using personal devices, typically in non-studio environments. This introduces variances including domestic sounds such as children, animals, and environmental noises. However, these real-world conditions can improve the robustness of the ASR system, helping it generalize better across diverse, noisy environments. While the recordings were done on single microphone devices, the platform used simulated stereo audio by duplicating the mono signal, resulting in a dual-mono format to ensure consistency. This approach does not introduce spatial depth or directionality, but it offers practical benefits: improved playback compatibility, cross-platform consistency, and the ability to later integrate true stereo recordings without restructuring the dataset.
@@ 4.7.1 Results
Figure 4: Image shows the reduction in common ASR errors when finetuned on the ViXSD data.
@@ 4.8 Lessons Learned
Transcriptions: Transcripts should be long form and conversational. This facilitates both easy quality control and a smoother workflow. Technical text or acronyms hinder the transcription processes as annotators often need to consult with the project managers to standardize how these are spoken.
@@
Recording Equipment Considerations: Participants employed various devices resulting in audio captured at different sample rates and stereo formats. While these differences required resampling and channel conversion, we observed no substantial quality gap between recordings. This suggests that strict device requirement may be unnecessary if proper recording guidelines are followed.
@@ 5 Conclusion
In this paper, we present the Esethu Framework, a pioneering approach to sustainable dataset governance that fundamentally reimagines how low-resource language data is created, licensed, and grown. Through its innovative reinvestment model—where licensing revenue is systematically channeled back into dataset expansion—the framework establishes a self-sustaining ecosystem for continuous resource development. As proof of concept, we developed the Vuk’uzenzele isiXhosa Speech Dataset (ViXSD), which demonstrates both the framework’s practical viability and its potential impact: our experiments show that fine-tuning yields up to 4.6% improvement in ASR performance, validating the technical quality of datasets produced under this approach. The Esethu Framework, with its emphasis on community ownership, sustainable growth, and ethical commercialization, offers a reproducible model for addressing the critical resource gaps in low-resource languages, ensuring that language communities remain primary beneficiaries of their data contributions while creating pathways for economic empowerment through language technology development.
@@ 6 Limitations
Given the ViXSD dataset focuses on news type articles, it may exhibit thematic and linguistic biases. Efforts have been made to preserve dialectal diversity and cultural authenticity, yet idiomatic expressions and culturally specific references may require careful interpretation in downstream NLP applications. Researchers should be cognisant of potential stereotypical or skewed linguistic representations when applying models trained on this corpus. Additionally, although efforts were made to ensure diverse speakers, the current dataset as it stands may not fully representative of the full socio-economic and geographical range of isiXhosa speakers. We also acknowledge that our current dataset remains relatively small, which may lead to models overfitting on specific speaker characteristics or acoustic environments. Future expansions using our proposed framework will include speakers from multiple provinces and regions to broaden coverage.
@@ 7 Ethical considerations
The creation of this ViXSD dataset necessitated careful consideration of ethical issues, particularly regarding data privacy, cultural representation, and linguistic bias. The dataset is derived from publicly available material, ensuring compliance with copyright and data protection regulations. Personally identifiable information (PII) is either absent or anonymised to safeguard individual privacy. Additionally, through the Esethu license, the data curators remain stewards of the dataset and any licensing fees for the data are reinvested into the development of the isiXhosa language speaker community, further data rights are detailed in Appendix A.5. The isiXhosa contributors were compensated at a fair market rate for their participation.
@@ Appendix A Appendix
A.1 Extracts from the Esethu License
@@ 3. We may permit commercial use of the dataset on the following terms:
A.2 Strengths and Weaknesses of the Esethu license
Table 5 provides an in-depth legal comparison of popular licenses with a focus on those developed for low-resource languages in the African context.
A.3 Extracts from the Commercial Agreement
@@
Metric Nwulite Obodo CC BY-SA 4.0 Apache 2.0 Kaitiakitanga Esethu License
Privacy No explicit protections. No provisions. No provisions. Implicit: Permission-based access. No explicit protections.
Agency for Data Creators High: Tiered control by community. Low: Licensor retains attribution rights. Moderate: Broad permissions granted. Highest: Māori custodians retain veto power. High: Co-owners control commercialisation.
Socioeconomic Benefits Royalties from non-developing countries can fund local data but not mandatory. None. None. Conditional: Commercial terms negotiable. Explicit: Commercial fees fund future African language datasets (clause 5.2).
Attribution Required. Required. Required (NOTICE file). Unspecified. Required under CC BY-SA for non-commercial use; commercial terms require attribution to data owners “where possible”.
Non-Commercial Use Free for Developing Countries; fees for others. Allowed. Allowed. Default: Prohibited unless permitted. Dual: Non-commercial = CC BY-SA; commercial = Esethu (fee-based).
No Modifications/Derivatives Modifications allowed (ShareAlike). Modifications allowed (ShareAlike). Modifications allowed (no ShareAlike). Restrictive: Derivatives bound by original license. Non-commercial = CC BY-SA (ShareAlike); commercial = Esethu (negotiated).
ShareAlike Required for derivatives. Required for derivatives. Not required. Strict: All derivatives bound by original license. Non-commercial = CC BY-SA (ShareAlike); commercial = Esethu (no ShareAlike).
Royalties Yes (non-developing countries). May include non-monetary benefit/interest. No. No. Potential (negotiated). Yes: Commercial fees paid to partner for reinvestment.
Open/Restrictive Hybrid: Geofenced openness. Fully Open. Permissive. Restrictive. Hybrid: Non-commercial = open (CC BY-SA); commercial = restrictive (Esethu).
Geographic Scope Tiered (Developing vs. non-Developing). Universal. Universal. Unspecified (prioritizes Indigenous users). Universal: Commercial terms apply globally, but fees waived for African entities (Annexure A, term 3).
Data Sovereignty Strong: Community reinvestment. Weak. Weak. Strongest: Embeds Indigenous governance. Strong: Co-owners retain control; proceeds fund African language datasets.
Enforcement of Derivatives Prohibits export to non-Developing Countries. Global sharing allowed (ShareAlike). No restrictions. Absolute: Derivatives must comply. Non-commercial = CC BY-SA terms; commercial = negotiated.
Termination Automatic if country reclassified. Violations only. Patent litigation only. Implicit (violations = revoked access). Breach triggers termination (clause 8).
Patent Rights Not addressed. Not addressed. Explicit grant. Not addressed. Not addressed.
Commercialization Control Permitted but non-developing nations subject to a royalty. No restrictions. No restrictions. Total control: Requires negotiation. Controlled: Commercial use requires fee (waived for African entities based on broad definition and case-by-case consideration).
License Compatibility Compatible with similar community licenses. Compatible with CC BY-SA. Compatible with most licenses. Incompatible: No relicensing allowed. Dual: CC BY-SA (non-commercial) + Esethu (commercial).
Table 5: Legal comparison of popular licenses with a focus on those developed for low-resource languages/the African context. We have used the metrics in the first column to compare them.
A.4 Speaker Guidelines
  1. 1.
    Recording Format
    1. (a)
      Record audio in WAV format.
    2. (b)
      Use free recording software (e.g., Audacity) as suggested.
  2. 2.
    Equipment Use
    1. (a)
      Use available recording devices: mobile phones, internal laptop microphones, or headset microphones.
  3. 3.
    Submission Process
    1. (a)
      Submit recordings via email or shared link as instructed.
  4. 4.
    Quality Assurance
    1. (a)
      Submit a sample recording for initial evaluation.
    2. (b)
      Feedback will be provided on background noise, reading fluency, and pacing.
    3. (c)
      Re-record prompts if there are misreads, hesitations, or unnatural pauses.
  5. 5.
    Language Use
    1. (a)
      Read numbers in isiXhosa rather than English.
  6. 6.
    Reading Guidelines
    1. (a)
      Read each prompt at a comfortable pace.
    2. (b)
      Record in a quiet room to avoid background noise.
  7. 7.
    Post-Processing
    1. (a)
      Recordings will undergo automated text–audio alignment and manual spot-checking.
    2. (b)
      Only clear, intelligible recordings closely matching the target text will be retained.
A.5 Data protection rights consideration.
Contributors assign their data to the creators, making the transfer irrevocable to ensure dataset continuity. However, contributors retain moral rights. Moreover, as voice recordings can be considered personal information, participants can invoke privacy laws (e.g., South Africa’s POPIA laws) to request limited data usage or corrections via a formal process.